OpenStudy (anonymous):

Malloy solved the equation −5x − 16 = 8; his work is shown below. Identify the error and where it was made. −5x − 16 = 8 Step 1: −5x − 16 + 16 = 8 + 16 Step 2: −5x = 24 Step 3: negative five times x over five = twenty-four over five Step 4: x = twenty-four over five Step 1: He should have subtracted 16 from each side. Step 1: He should have divided both sides by −5. Step 3: He should have multiplied both sides by −5. Step 3: He should have divided both sides by −5.

OpenStudy (anonymous):

he should have divided both sides by -5

OpenStudy (compassionate):

The steps to solve the equation are: \[-5x - 16 = 8\]\[-5x - 16 + 16 = 8 + 16 \]\[-5x = 24 \rightarrow \frac{-5x}{-5} = \frac{24}{-5} \]\[x = -4.8\]

OpenStudy (anonymous):

Step 3: He should have multiplied both sides by −5. Because in step 3 he divided both side by 5, however in the next step he ended up with positive x.